OpenStudy (anonymous):

4c + 14 = −3c + 8 c=

OpenStudy (anonymous):

this is a nail bitter

OpenStudy (stormfire1):

4c + 3c = 8 -14 7c = -6 c = -6/7

OpenStudy (lilg132):

4c + 14 = -3c + 8 4c + 3c = 8 - 14 7c = -6 c = -0.8571428571428571

OpenStudy (anonymous):

thank you soooo much

OpenStudy (stormfire1):

I'd go with -6/7 ...it's cleaner :)

OpenStudy (lilg132):

it is but just to be precise thats the answer :p

OpenStudy (lilg132):

and your welcome

OpenStudy (stormfire1):

-6/7 is more precise :)

OpenStudy (stormfire1):

For the sake of elaborating it for ilovethis21: -6/7 is more precise (and shorter) because the actual answer goes on for a lot more decimal places than that. -6/7 is considered an exact answer whereby the decimal answer is considered only an approximation.